OpenStudy (anonymous):

Given F(x) shown below, complete the equation for the inverse of F(x). F(x)= 2x/7 + 4 F-1(y)= ?

OpenStudy (anonymous):

solve it for x (let's call F(x)=y): y-4=2x/7 2x=7y-28 x=7y/2-14 now re-write is as inverse F-1(x): F-1(x)=7x/2 -14